Shift report 20010806 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Mon Aug 06 2001 - 16:08:59 EDT

  • Next message: Hironori Ito: "RE: New Si and BB dN/dEta at 100 Gev/A !!!"

    ------------------------------------------------------------
    Date: Mon Aug  6 16:08:58 2001, Shift: 08:00-16:00
    Supervisor: PC, with: 
    Log book#7 pages 34 - 36
    
    Shift Summary:
    ------------------------------------------------------------
    1) 
    End of night shift run at 9:45
    RF testing the rest of the day. Access at 20:00
    
    2) 
    Run 4002 ~240K collisions stopped when I noticed that T2
    had tripped.
    Run4005 (10 minutes ~ 5k collisions) with T1 and T2 off. 
    
    3) 
    Now both T1 and Anode A on T2 have problems. We will have
    an access at 20 today.
    
    Hopefully better news from the evening guys.
    
    The plan for data taking is :
    MRS at 90 2kG
    Change polarities between stores if have > 10^6 collisions
    FFS at 6 deg. 1/3 field B polarity till T1 and T2 fixed.
    
     
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Mon Aug 06 2001 - 16:09:22 EDT